OpenStudy (anonymous):

Perimeter of a rectangle. The width of a rectangle is three fourths of the length. The perimeter of the rectangle becomes 50 cm when the length and the width are each increased by 2 cm. Find the length and the width.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

let x and y be length and width of the rectangel. y = 3/4 x when x and y are increased by 2cm perimeter becomes 50 cm 50= 2 ( (x+2) + ( 3/4x +2) ) 50 = 2 ( 7x/2 + 4) 50 = 7x + 8 7x = 42 x=6 length, width = 9/2 cm

OpenStudy (anonymous):

thanks :)

OpenStudy (radar):

Perimeter would equal 2L + 2W then P= 2*6 + 2*( 9/2) P=12+9=21 Shouldn't it be 50? What am I missing or is the solution wrong!

OpenStudy (anonymous):

no idea lol copied it exactly out of my online book

OpenStudy (anonymous):

hey radar could you try my other problem?

OpenStudy (anonymous):

hey guys, perimeter will be 50cm when existing length and width are increased by 2 cm each

OpenStudy (anonymous):

nagendra, could u look at my other one?

OpenStudy (anonymous):

I have to go to work soon and Im stumped

OpenStudy (anonymous):

where is the other one

OpenStudy (anonymous):

starts with womens records...

OpenStudy (anonymous):

8th one down

OpenStudy (anonymous):

i'm sorry i made a small error in the previous one 50 = 2( 7x/4 +4 ) 50 = 7x/2 +8 7x/2 = 42 x= 12, y = 9 please note the corrections

OpenStudy (anonymous):

ok thanks :) could you check the other one out?

OpenStudy (anonymous):

please copy it here

OpenStudy (anonymous):

Women’s records in the women’s 1500-m run. The formula R =−0.0026? + 4.0807 can be used to predict the world record, in minutes, for the 1500-m run t years after 1900. Determine (in terms of an inequality) those years for which the world record will be less than 3.8 min.

OpenStudy (anonymous):

i posted it in the original post, is it ? or t

OpenStudy (radar):

Thanks for info. It does check out now. I've been outside retrieving a boat that washed over the spillway. P=2(12+2)+2(9+2)=28+22=50